4x+y=70 when 3x=y x=quarters and y=dollar bills so what is x and y equal to add up to $70 while still solving 3x=y
Found 2 solutions by oscargut, stanbon:
Answer by oscargut(2103)

Since 3x=y
4x+y = 4x+3x = 7x = 70
Then x = 10 and y = 30
If you need more help you can contact me at
mthman@gmail.com
Thanks

Answer by stanbon(75887)

4x+y=70 when 3x=y x=quarters and y=dollar bills so what is x and y equal to add up to $70 while still solving 3x=y
-------
Substitute for "y" and solve for "x"
4x + 3x = 70
7x = 70
x = 10
----
y = 3x = 30
===================
Cheers,
Stan H.
=============